Solve the following differential equation. Do not use Laplace. y – 4y = 2e (2x+3) - Write the corresponding homogeneous eq
Answer #1

y- 4y = 2622+3* Homogeneous egr = yl_qy=0 03-9074 0 - Amarillany Egnation in m3-gmzo 72) N 4 ( mº-4) 3 6 -DM20,+2=2 Then th

If you satisfied with this answer please give positive ratings.

Thank you.